# ZebraLogic: On the Scaling Limits of LLMs for Logical Reasoning

### v1

@@ Abstract
We investigate the logical reasoning capabilities of large language models (LLMs) and their scalability in complex non-monotonic reasoning. To this end, we introduce ZebraLogic, a comprehensive evaluation framework for assessing LLM reasoning performance on logic grid puzzles derived from constraint satisfaction problems (CSPs). ZebraLogic enables the generation of puzzles with controllable and quantifiable complexity, facilitating a systematic study of the scaling limits of models such as Llama, o1 models, and DeepSeek-R1. By encompassing a broad range of search space complexities and diverse logical constraints, ZebraLogic provides a structured environment to evaluate reasoning under increasing difficulty.
Our results reveal a significant decline in accuracy as problem complexity grows—a phenomenon we term the “curse of complexity.” This limitation persists even with larger models and increased inference-time computation, suggesting inherent constraints in current LLM reasoning capabilities. Additionally, we explore strategies to enhance logical reasoning, including Best-of-N sampling, backtracking mechanisms, and self-verification prompts. Our findings offer critical insights into the scalability of LLM reasoning, highlight fundamental limitations, and outline potential directions for improvement.
@@
byuchen@uw.edu  ronanlb@allenai.org  yejinc@stanford.edu https://hf.co/spaces/WildEval/ZebraLogic
@@ 2.2 Problem Formulation
Clue-Based Constraints: Each clue in the puzzle introduces additional constraints that must be satisfied by any valid assignment [^2] Note that there are also several implicit positional constraints that must be considered. For example, the leftmost house cannot be on the right of any other house, and the rightmost house cannot be on the left of any other house (as relevant in Clue 4). These spatial constraints, combined with the explicit clues, translate the verbal descriptions into precise logical conditions to be satisfied.. Under the hood, these clues are translated into formal logic formulas that constrain the relationships between variables. For our example puzzle in Figure 2, the constraints can be formulated as follows:
@@ 6.2 Self-Refinement is Limited but Promising
The other feature of o1’s hidden reasoning process is the ability to reflect on its own reasoning process and refine its answer. From our observation on the summary of their hidden reasoning process, we can see that o1 often revisits the clues and constraints to verify its previous reasoning and fix the errors if there are any, which is similar to the Z3 solver’s conflict-driven clause learning mechanism. In order to elicit such self-refinement behavior from LLMs, we add follow-up queries to ask the model to review its initial answer and check the clues and constraints again in a multi-turn conversation setting. There are two settings for the self-refinement process: one with the oracle knowledge of the correct answer and the other without the oracle knowledge. Results in Table 2 show modest improvements with self-verification, particularly without oracle knowledge (4o improves from 31.7 to 33.0, then decreases to 32.1 on second iteration).
@@ 7 Related Work
Approaches to Logical Reasoning in LLMs
Several lines of research propose methods to augment or refine LLMs for stronger logical reasoning. Clark et al. (2020) demonstrated that transformers can emulate logical reasoning over natural language sentences—serving as “soft theorem provers.” Pan et al. (2024) showed that a decoder-only Transformer could tackle SAT problems, paralleling the Davis–Putnam–Logemann–Loveland (DPLL) algorithm, thereby expanding the role of LLMs to more complex problem-solving domains. Alternatively, neuro-symbolic systems like CLOVER (Ryu et al., 2024) integrate LLMs with symbolic solvers to better capture the translation of intricate logical semantics from text.
Empirical Evidence of LLM Limitations
Despite these promising developments, LLMs face persistent hurdles as logical problem complexity increases. Yan et al. (2024) contended that models may rely heavily on probabilistic correlations rather than genuinely understanding logical rules. Similarly, Xie et al. (2024) highlighted the complex interplay between training data memorization and genuine reasoning abilities of LLMs. Additionally, Schlegel et al. (2022) conducted an extensive empirical study to investigate the detection of formally valid inferences in controlled fragments of natural language, revealing that transformers often overfit to superficial patterns rather than acquiring logical principles. Lam et al. (2024) showed the impact of the choice of symbolic solvers on the effectiveness of LLMs in deductive reasoning tasks, calling for more consistent comparative studies. Further empirical evidence from Dziri et al. (2023) and Parmar et al. (2024) demonstrated that even ostensibly simple logical tasks continue to challenge these models. Finally, Madusanka et al. (2023) investigated the limits of transformers on solving the problem of model-checking with natural language and the significant impact of the language fragment on the performance of transformers.
@@ 8 Conclusion
We introduce ZebraLogic, a controlled benchmark of 1,000 logic grid puzzles that highlights the scaling limits of LLM-based reasoning through carefully adjustable complexity. Our experiments reveal a pronounced drop in model performance as puzzle complexity increases, overshadowing gains from model growth or training data expansions. While increasing the generation sample size yields modest improvements, a backtracking-based approach with expanded reasoning steps significantly boosts accuracy. These results spotlight the importance of explicit, step-by-step reasoning strategies and provide a valuable framework for advancing logical reasoning research in LLMs.
@@ Appendix D Further Discussion on o1’s Reasoning
Visible outputs from o1 cannot fully explain its reasoning for complex problems. To understand how o1 reasons, we have to focus on their public reasoning steps that we can extract from the model’s visible outputs. From our human evaluation on their reasoning steps, we find that o1’s reasoning steps are not necessarily rigorous or complete, even when they arrive at the correct solution. For small-to-medium search spaces, o1-preview’s reasoning chains tend to be complete, while o1-mini sometimes can skip some steps to directly reach the solution. For problems with larger search spaces, o1’s visible reasoning chains tend to be very incomplete, and sometimes even incorrect, especially when the reasoning process requires backtracking. For example, o1’s visible reasoning may contain steps such as “Bob cannot be in Houses 1, 4, or 5, so he must be in House 3” without explaining why Bob cannot be in House 2, although it will indeed lead to the correct solution. Note that such cases also happen for other LLMs such as GPT-4o. We thus describe that the reasoning process of LLMs and o1 models are sometimes based on guessing without formal logic, especially for complex problems with large search spaces, rater than rigorous logical reasoning.

### v2

@@ Abstract
We investigate the logical reasoning capabilities of large language models (LLMs) and their scalability in complex non-monotonic reasoning. To this end, we introduce ZebraLogic, a comprehensive evaluation framework for assessing LLM reasoning performance on logic grid puzzles derived from constraint satisfaction problems (CSPs). ZebraLogic enables the generation of puzzles with controllable and quantifiable complexity, facilitating a systematic study of the scaling limits of models such as Llama, o1 models, and DeepSeek-R1. By encompassing a broad range of search space complexities and diverse logical constraints, ZebraLogic provides a structured environment to evaluate reasoning under increasing difficulty. Our results reveal a significant decline in accuracy as problem complexity grows—a phenomenon we term the “curse of complexity.” This limitation persists even with larger models and increased inference-time computation, suggesting inherent constraints in current LLM reasoning capabilities. Additionally, we explore strategies to enhance logical reasoning, including Best-of-N sampling, backtracking mechanisms, and self-verification prompts. Our findings offer critical insights into the scalability of LLM reasoning, highlight fundamental limitations, and outline potential directions for improvement.
@@
byuchen@uw.edu  ronanlb@allenai.org  yejinc@stanford.edu https://hf.co/spaces/allenai/ZebraLogic
@@ 2.2 Problem Formulation
Clue-Based Constraints: Each clue in the puzzle introduces additional constraints that must be satisfied by any valid assignment. Note that there are also several implicit positional constraints that must be considered. For example, the leftmost house cannot be on the right of any other house, and the rightmost house cannot be on the left of any other house (as relevant in Clue 4). These spatial constraints, combined with the explicit clues, translate the verbal descriptions into precise logical conditions to be satisfied.. Under the hood, these clues are translated into formal logic formulas that constrain the relationships between variables. For our example puzzle in Figure 2, the constraints can be formulated as follows:
@@ 6.2 Self-Refinement is Limited but Promising
The other feature of o1’s hidden reasoning process is the ability to reflect on its own reasoning process and refine its answer. From our observation on the summary of their hidden reasoning process, we can see that o1 often revisits the clues and constraints to verify its previous reasoning and fix the errors if there are any, which is similar to the Z3 solver’s conflict-driven clause learning mechanism. In order to elicit such self-refinement behavior from LLMs, we add follow-up queries to ask the model to review its initial answer and check the clues and constraints again in a multi-turn conversation setting. There are two settings for the self-refinement process: one with the oracle knowledge of the correct answer and the other without the oracle knowledge. Results in Table 2 show modest improvements with self-verification, particularly without oracle knowledge (4o improves from 31.7 to 33.0, then decreases to 32.1).
@@ 7 Related Work
Approaches to Logical Reasoning in LLMs. Several lines of research propose methods to augment or refine LLMs for stronger logical reasoning. Clark et al. (2020) demonstrated that transformers can emulate logical reasoning over natural language sentences—serving as “soft theorem provers.” Pan et al. (2024) showed that a decoder-only Transformer could tackle SAT problems, paralleling the Davis–Putnam–Logemann–Loveland (DPLL) algorithm, thereby expanding the role of LLMs to more complex problem-solving domains. Alternatively, neuro-symbolic systems like CLOVER (Ryu et al., 2024) integrate LLMs with symbolic solvers to better capture the translation of intricate logical semantics from text.
Empirical Evidence of LLM Limitations. Despite these promising developments, LLMs face persistent hurdles as logical problem complexity increases. Yan et al. (2024) contended that models may rely heavily on probabilistic correlations rather than genuinely understanding logical rules. Similarly, Xie et al. (2024) highlighted the complex interplay between training data memorization and genuine reasoning abilities of LLMs. Additionally, Schlegel et al. (2022) conducted an extensive empirical study to investigate the detection of formally valid inferences in controlled fragments of natural language, revealing that transformers often overfit to superficial patterns rather than acquiring logical principles. Lam et al. (2024) showed the impact of the choice of symbolic solvers on the effectiveness of LLMs in deductive reasoning tasks, calling for more consistent comparative studies. Further empirical evidence from Dziri et al. (2023) and Parmar et al. (2024) demonstrated that even ostensibly simple logical tasks continue to challenge these models. Finally, Madusanka et al. (2023) investigated the limits of transformers on solving the problem of model-checking with natural language and the significant impact of the language fragment on the performance of transformers.
@@ 8 Conclusion
We introduce ZebraLogic, a controlled benchmark of logic grid puzzles that highlights the scaling limits of LLM-based reasoning through carefully adjustable complexity. Our experiments reveal a pronounced drop in performance as complexity increases, overshadowing gains from model growth or training data expansions. While increasing the generation sample size yields modest improvements, a backtracking-based approach with expanded reasoning steps significantly boosts accuracy. These results spotlight the importance of non-monotonic reasoning and provide a valuable framework for advancing logical reasoning research.
Acknowledgments
Yejin Choi’s research is supported in part by the National Science Foundation under Grant DMS-2134012.
@@ Appendix D Further Discussion on o1’s Reasoning
Visible outputs from o1 cannot fully explain its reasoning for complex problems. To understand how o1 reasons, we have to focus on their public reasoning steps that we can extract from the model’s visible outputs. From our human evaluation on their reasoning steps, we find that o1’s reasoning steps are not necessarily rigorous or complete, even when they arrive at the correct solution. For small-to-medium search spaces, o1-preview’s reasoning chains tend to be complete, while o1-mini sometimes can skip some steps to directly reach the solution. For problems with larger search spaces, o1’s visible reasoning chains tend to be very incomplete, and sometimes even incorrect, especially when the reasoning process requires backtracking. For example, o1’s visible reasoning may contain steps such as “Bob cannot be in Houses 1, 4, or 5, so he must be in House 3” without explaining why Bob cannot be in House 2, although it will indeed lead to the correct solution. Note that such cases also happen for other LLMs such as GPT-4o. We thus describe that the reasoning process of LLMs and o1 models are sometimes based on guessing without formal logic, especially for complex problems with large search spaces, rather than rigorous logical reasoning.